What is the frequency of light (in Hz) of a photon emitted when an electron in a hydrogen atom undergoes a transition from the n = 6 energy state to the n = 2 energy state? Express your answer to two significant figures.
1
Expert's answer
2018-10-26T08:23:09-0400

What is the frequency of light (in Hz) of a photon emitted when an electron in a hydrogen atom undergoes a transition from the n = 6 energy state to the n = 2 energy state? Express your answer to two significant figures.

n1= 6

n2= 2

Z= 1

R= 3.29x10^15 c^-1

nu--?

Solution

nu = R*(1/(n2)^2 - 1/(n1)^2)= 3.29x10^15*(1/4 - 1/36)=0.73 x10^15= 7.3x10^14c-1

Answer: 7.3x10^14c-1